OpenStudy (anonymous):

Estimate √55 to the nearest whole number.

OpenStudy (dls):

\[\LARGE f(x+ \Delta x)= f(x)+\Delta x(f'(x))\] Let \(\Large f(x)=\sqrt{x}\) Replace x with 7 and \(\Delta x\) with 6

OpenStudy (nirmalnema):

7.41

OpenStudy (nirmalnema):

u can say it 7 because it is < 7.5

OpenStudy (anonymous):

thank you :)

OpenStudy (dls):

or \[\LARGE 7 \times 7=49\] \[\LARGE 8 \times 8 =64\] \[\LARGE \frac{64+49}{2}=56.5\] so it should be the average between 7 and 8 \[\LARGE \frac{7+8}{2}=7.5\] is the answer